The daughter is 3 years old and her mother is 28 older than her daughter. how many years later will be the mother be 3times old as her daughter

Answer by ankor@dixie-net.com(22740)

The daughter is 3 years old and her mother is 28 older than her daughter.
mom is 31
how many years later will be the mother be 3times old as her daughter
let y = no. of yrs for this to be true
31+y = 3(3+y)
31 + y = 9 + 3y
31 - 9 = 3y - y
22 = 2y
y = 11 years for this to be true
:
:
Check this, find their ages in 11 yrs
Mom; 31+11 = 42
Daughter will be 3+11 = 14
Mom 3 times at old as the daughter